OpenStudy (anonymous):

When planning his crops, Farmer Julie knows that her 25 acres can support oranges and pears. She wants to make $7,500 from her crops. She can make $2,050 per acre of oranges (the variable o) and $2,000 per acre of pears (the variable p). Which equation below would be a constraint in her system of equations?

OpenStudy (anonymous):

Is o + p <= 25 one of the options? Cannot plant more acres than you have.

OpenStudy (anonymous):

2,050o - 2,000p = 7,500 2,050o + 2,000p = 7,500 o - p = 25 o + p = 25 Sorry went afk so C) ot D)

OpenStudy (anonymous):

o + p = 25 right?

OpenStudy (anonymous):

Yes, o + p = 25, assuming he uses all his acreage.